OpenStudy (anonymous):

Ace Electronics, Hi-Fi Stereo, and CD World buy a total of 170 VCR/DVD players from the same distributor each month. Because CD World buys 30 more units than the other two stores combined, CDW's cost is only $160 per unit. The players cost Hi-Fi $165 each and Ace $170 each. How many players does each retailer buy each month if the distributor receives $27,675 each month from the sale of the players to the three stores? Ace Electronics units Hi-Fi Stereo units CD World units

OpenStudy (anonymous):

a+h+c=170 c=30+h+a 160c+165h+170a=27675 a=25,c=100,h=45

OpenStudy (anonymous):

Thank you so much